# SARATR-X: A Foundation Model for Synthetic Aperture Radar Images Target Recognition

### v1

@@ 4.3 Design of target signal for SAR images
First, we consider whether the existing methods based on the ViT are suitable for SAR images. PixMIM [46] applies the low pass filter to remove the high-frequency components and drives the model to focus on shape information. However, PixMIM does not outperform MAE because the noise type of SAR is multiplicative, and the filter parameters require a trade-off between target and noise. Then, FG-MAE [79] uses HOG to capture the feature for the SSL with SAR scene-level tasks, but we find that HOG does not ensure high-quality SAR target features. Target regions usually have strong scattering values, and the speckle noise causes the gradient computation to have strong false points in these regions. Besides, I-JEPA [2] propose deep networks as target feature encoders to capture the deep semantic features but lead to training overfitting noise and failure to learn effective feature representations.
@@
Therefore, we choose the SAR features as target features to enhance HiViT. SAR-HOG changes the gradient calculation of HOG features and uses the gradient by ratio to solve speckle noise, and its result performs better than pixel value and HOG. Inspired by PixMIM, we prefer to use the target shape (i.e., gradient features) directly as the target feature  [^2] SAR-HOG uses the same multi-scale setting to illustrate that simple gradient features can effectively represent target shape as an SSL guide signal. Besides, multi-scale can improve the feature representation for various small targets in remote sensing. We discuss kernel settings for computing gradient in Fig. 4. Scale can affect feature quality: a smaller scale is finer for small target edge extraction, while a larger scale is more suitable for large target and noise suppression. Therefore, combining features of different scales has an improvement over a single scale for various target sizes in images.
Figure 4: Kernel settings for computing gradient by ratio. Here, the scale 1/2/3 is with $r$ equal to 9/13/17, and the multi-scale contacts all scales. Multi-scale is more suitable than single scale for various targets in remote sensing images.
4.4 Analysis
@@
As stated above, SARATR-X’s key points are summarized as follows: HiViT architecture avoids the loss of small target information. SSL-ImageNet & SAR use ImageNet pre-training weights to provide a good initialization for diversity perceptual capability. MGF ensures high-quality target features and suppresses speckle noise under SSL with SAR images. Benefiting the above insights, our SARATR-X can learn high-quality target features from noisy SAR remote sensing images in Table IV. Next, we analyze the diversity and scalability of SARATR-X.
@@
Figure 5: Averaged attention distance in different attention heads (x-axis is the attention head w.r.t layer number) on SSL models. Attention distance represents the range of the receptive field. We can focus on model architectures (Fig. (a) v.s. Fig. (b)), initialization weights (Fig. (a) v.s. Fig. (c)), and SSL signals (Fig. (d) v.s. Fig. (e)) to ensure diverse attention ranges in SAR target recognition, such as HiViT architecture, ImageNet weights, and SAR target features.
@@
Figure 6: Scaling ability of SARATR-X in dataset size, parameters, and training epochs with linear probing performance. Our method can benefit from these three aspects. However, it is important to note that excessive training epochs can lead to overfitting due to the dataset size and speckle noise.
@@
Visualization. Research [85] has shown that supervised pre-training and contrastive learning only model global information in the high layer, and MIM can model both local and global information. However, we observe that this phenomenon is not only related to the methods but also to the data properties. Fig. 5 (a) shows that the ViT with MAE focuses on global information due to large SAR image scenes, which differs from MIM’s modeling properties. Therefore, HiViT has various attention ranges with its high spatial resolution and hierarchical structure in Fig. 5 (b). In addition, using ImageNet’s weight as initialization can enhance this problem in Fig. 5 (c). As for the target feature, Fig. 5 (d) displays the HOG feature enhances the noise interference, which harms feature diversity. MGF effectively extracts the shape information of the target, making the model more focused on diverse edge information in the lower layer. However, this approach removes texture and preserves edge information, which motivates the higher layers not to need to focus on texture details and diminishes the attention range in Fig. 5 (e). Therefore, we combine two-step training with MGF in Fig. 5 (f).
@@
Scaling experiment. Although MIM is a good learner that scales with data and model resources [86], a question arises as to whether our method can ensure scalability for MIM when dealing with noisy data such as SAR. Fig. 6 presents the scaling experiment from three perspectives: dataset size, parameters, and training epochs. Despite our pre-training set comprising 186,660 images, which is smaller than ImageNet-1K, we observe a significant improvement in downstream tasks’ performance with increasing data and parameters. This phenomenon indicates that the foundational model can fully unleash its potential in SAR images by extracting high-quality features as guiding signals. However, just like the study [86], when the pre-training size contains 100,000 images, the model tends to overfit during extended training epochs. Moreover, SAR image noise and low resolution can further aggravate the overfitting. However, SARATR-X outperforms our previous study, SAR-JEPA, which has overfitted at 400 epochs with 94,776 SAR images. There is a need to continue investigating how to ensure high-quality feature representation when extending the SAR foundation models.
5 Leveraging SARATR-X for Recognition
@@
We have discussed different aspects of SARATR-X, but there are many specific datasets and specialized models for SAR ATR. We thus compare our SARATR-X with other state-of-the-art methods, such as supervised learning ($\rm{CS}^{n}$et [9] and PD [95]), semi-supervised learning (EUAPS [100]), and self-supervised learning (MSFA [43] and BIDFC [92]). We focus on SAR recognition tasks, including image classification and detection. More detailed settings [^3] We removed downstream tasks’ test sets from the pre-training set samples. are in appendix B.
TABLE VI: SAR classification results on MSTAR’s SOCs and EOCs settings. SOCs are in a similar distribution and have ten types of vehicle targets. EOCs include variations for imaging conditions that belong to robustness classification and out-of-distribution problems. EOCs use depression angle, configuration, and version variations to test robustness. We find that self-supervised (BIDFC and Ours) and semi-supervised (EUAPS) methods significantly perform under the few-shot task with additional unlabeled data, demonstrating the data’s importance. Besides, our method is robust to variations in imaging parameters. The metric is $\uparrow$, and the detail setting follows the $N$-shot [94]. Bold indicates the best result, underline is the next best result. We provide detailed results in Appendix B.
@@
SARATR-X 2024 65.3 (+11.8) 76.5 (+20.3) 82.8 (+15.6)
TABLE VII: SAR detection results on SARDet-100K, OGSOD, SSDD, and SAR-Aircraft. Our proposed SARATR-X has competitive performance on various detection datasets. Bold indicates the best result, underline is the next best result. More detailed results are in Appendix B.
@@
Classification task. We study in Table VI the performance of SARATR-X on the MSTAR [1] dataset with Standard Operating Conditions (SOCs) and Extended Operating Conditions (EOCs). We first notice that SSL (BIDFC and ours) and semi-supervised (EUAPS) significantly outperform other methods under small samples with additional unlabeled samples. Our results surpass the previous best by large margins, demonstrating the value of foundation models in an era of rapidly growing SAR data. In particular, our SARATR-X shows robustness to the EOC setting of imaging condition variations. This result indicates that the foundation model learns stable features and relationships from diverse imaging conditions in a large number of target samples.
Detection task. As illustrated in Table VII, we report box AP for SAR target detection with a horizontal bounding box on multi-category (SARDet-100K and OGSOD), ship (SSDD), and aircraft detection (SAR-Aircraft). SARATR-X performs better than our previous MSFA by 0.8 points on SARDet-100K. MSFA has more complex training processes and target features, which use multi-stage training between RGB and SAR images with three different target features and have a detection pre-training step. SARATR-X is more simple but more effective for SAR images. More significantly, SARATR-X outperforms or has comparable performance on various datasets compared to many specifically designed detection methods in Table VII. Of course, our study is only a preliminary exploration of SSL for SAR. More effective target features can be realized in a data-knowledge dual-driven manner by further mining the knowledge of SAR imaging mechanisms and properties. Moreover, with the larger dataset and parameters, the path of foundation models will hopefully lead to general SAR target recognition, but this goal requires researchers together.
6 Conclusion and Future Perspectives
In this paper, we proposed the SARATR-X for SAR ATR and systematically investigated a foundation model framework. First, a pre-training dataset was built from 14 open-source datasets, including various targets, scenes, and sensors. Then, the foundation model’s pre-training backbone, SSL methods, and downsteam tasks were discussed in detail. Importantly, SARATR-X demonstrated superior performance on different target recognition datasets, demonstrating the foundation model’s potential in this field.
@@ 6 Conclusion and Future Perspectives
We believe that research on SAR foundation models, such as SARATR-X, has the potential for generalized feature representations in SAR images and accelerates towards all-day, all-weather target recognition in Earth observation. However, the research of the foundation model requires large data, which is a real problem for SAR images. SAR images are expensive and require specific imaging equipment and algorithms, and privacy and security prevent the data from opening. Therefore, we are particularly grateful to the publishers of open-source SAR target datasets. By making SARATR-X publicly available, we aim to accelerate the progress of the foundation model in SAR target recognition by enabling researchers to use our dataset and code to design better methods or explore downstream applications.
@@
Although this work systematically investigated a foundation model framework, several limitations and challenges require exploration in future work. The SAR images are derived from open-source SAR datasets, and the targets are mainly vehicles, ships, aircraft, oil tanks, etc. Collecting target samples from increasingly unlabelled SAR imagery could further expand the amount of data and the range of downstream applications. In addition, investigating the expert knowledge with text for multimodal interactions to describe the relationship between targets and scenes can further enhance the representation capability of the foundation model. Due to the visual variability of SAR images, textual descriptions are an effective tool for improving interpretability and comprehensibility.
@@
In conclusion, we have verified the SARATR-X’s ability to adapt to diverse SAR target datasets, showing high performance and generalizability in classification and detection. By taking full advantage of the rapid growth of SAR images, the SSL-based foundation model opens the door to a generalized feature representation and SAR target recognition. We believe it’s time to embrace fundamental models for SAR image interpretation.
Appendix A Implementation Details of Section 4 Experiments of SARATR-X
@@
A.1 Pre-traing dataset setting
@@
As shown in Fig. 7, we collect data from open-source datasets based on our previous research [38, 43]. Now, our pre-training dataset contains 14 open-source SAR target datasets. Here are brief descriptions of each dataset’s targets, scenes, and sensors.
@@
A.2 Classification dataset for performance Test
TABLE VIII: Description of our SAR classification dataset, named SAR-VSA, which contains 25 fine-grained targets. # Train: Number of train samples. # Test: Number of test sample images.
@@ A.2 Classification dataset for performance Test
We select three target classification datasets, including 25 fine-grained targets from vehicles, ships, aircraft, and others, to evaluate the comprehensive performance of SSL and the foundation model for SAR target recognition. The new SAR classification dataset named SAR-Target in Table VIII.
A.3 Hyperparameter settings
Here are the detailed settings of our pre-training and downstream tasks as shown in Table IX and X.
Pre-training. Our default pre-training setting is in Table IX, and other hyperparameters of each method use the default settings from their papers and codes. Our pre-training is applied on 8 NVIDIA RTX3090 GPUs with 200 epochs and 800 batch sizes. Compared to the training settings of MAE, we add ColorJitter ($\rm{contrast}=0.5$) to increase data richness. Moreover, we modify the batch size and epoch according to 8 GPUs. It is worth noting that although MAE uses the normalized pixel value to enhance the feature representation in the visible spectral images, we find that the normalized pixel value cannot be used due to the SAR image noise and prevents the training loss from decreasing properly.
TABLE IX: Pre-training setting.
@@ A.3 Hyperparameter settings
Classification setting. All models use the same training settings in downstream classification tasks. Table X gives the default setting. Our few-shot learning setting is based on the Dassl toolbox [101, 102] and is averaged over 10 random experiments. Since we focus on the small sample case of the downstream classification task, we use the linear probing method in MAE to finetune models and avoid overfitting.
TABLE X: Classification settings (fine-tuning and linear probing).
@@
Partial fine-tuning. Fig. 8 shows why we chose linear probing for few-shot evaluation, as HiViT overfits when many blocks are fine-tuned. Experimental results show that our SAR-KGPA consistently obtains better representations than MAE, and overfitting occurs later.
Appendix B Implementation Details of Section 5 Leveraging SARATR-X for Recognition
B.1 Dataset description
We choose MSTAR, the most commonly used dataset in SAR target classification. SOCs are in similar image conditions, and the training set’s depression angle under SOC is 17°, while the test set is 15°. Ten categories of targets include BMP2, BRDM2, BTR60, BTR70, T62, T72, 2S1, D7, ZIL131, and ZSU234 as shown in Table VIII. The EOCs setting is the imaging condition variations to test robustness. Existing methods have been saturated with different experimental settings of MSTAR [40], but few samples and depression angle variations remain challenging. We use SARDet-100K and OGSOD datasets, which have many test samples and categories, to evaluate the detection performance fully. The OGSOD comparison results are derived from the original article’s [71] single-modal approach using only SAR images. SSDD and SAR-Aircraft are ship and aircraft categories.
B.2 Hyperparameter settings
Classification setting is follow Table X. The only difference is that we use partial fine-tuning for better performance, and the last 6 blocks are added to the fine-tuned.
B.3 Detailed results
We provide detail classification and detection result on Table XI and XII. Although the proposed method outperforms existing methods in mAP, there is still scope for improvement in some refined metrics.
TABLE XI: Detailed classification results on MSTAR’s SOCs and EOCs. Bold indicates the best result, underline is the next best result.
@@ B.3 Detailed results
SARATR-X 2024 65.3 76.5 76.3 83.4 82.8 85.0 92.1 97.0
TABLE XII: Detailed detection results. The metrics are mAP, $\rm{mAP_{50}}$ (@@@@50), $\rm{mAP_{75}}$ (@@@@75), $\rm{mAP_{small}}$ (@@@@s), $\rm{mAP_{medium}}$ (@@@@m), and $\rm{mAP_{large}}$ (@@@@l). Bold indicates the best result, underline is the next best result.
@@
SARATR-X 2024 58.7 86.1 (+5.7) 64.7 (+3.3) 20.0 57.2 49.7

### v2

@@ IV-C Design of target signals for SAR images
After investigating the backbone and strategy, we focused on target features for SSL methods used with SAR images. One key point for MIM is designing high-quality guide signals due to the unique multiplicative speckle noise in SAR images. This means we need to suppress noise and enhance target features. As seen in Table III, we considered five target features (pixel values) [81], including a low pass filter [90], HOG features [43], deep features [91], SAR-HOG [85], and gradient by ratio [86, 85]). All SSL methods used the SSL-SAR setting and base version. We first considered whether existing methods based on ViT were suitable for SAR image classification. PixMIM [90] applies a low pass filter to remove high-frequency components, driving the model to focus on shape information. However, PixMIM did not outperform MAE because the noise type in SAR is multiplicative, and the filter parameters require a trade-off between the target and noise. FG-MAE [43] uses HOG to capture SSL features in SAR scene-level tasks, though we found that HOG did not ensure accurate SAR target features. Target regions typically exhibited strong scattering values, and the included speckle noise often caused the gradient computations to exhibit strong false points in these regions. In addition, I-JEPA [91] proposed deep networks for use as target feature encoders to capture deep semantic features. However, this can lead to training overfitting noise and a failure to learn effective feature representations.
@@
As such, we chose SAR features as target features to enhance HiViT. SAR-HOG changes the gradient calculations for HOG features and uses gradient by ratio to solve for the speckle noise, thereby outperforming pixel values and HOG. Inspired by PixMIM, we prefer to directly use the target shape (i.e., gradient features) as a target feature  [^3] SAR-HOG uses the same multi-scale settings to illustrate that simple gradient features can effectively represent target shape as an SSL guide signal.. In addition, multi-scale methods can improve the feature representations of various small targets common in remote sensing. Discussions of kernel settings used for computing gradients are illustrated in Fig. 3. Scale can also affect feature quality, as a smaller scale is finer for small target edge extraction, while a larger scale is more suitable for large targets and noise suppression. Therefore, combining features of different scales (see Fig. 3) offers improvements over a single scale for various image target sizes.
IV-D Analysis
@@
As stated above, the primary contributions of SARatrX can be summarized as follows: the HiViT architecture avoids the loss of small target information. SSL-ImageNet & SAR use ImageNet pre-training weights to provide a good initialization for diversity perceptual capabilities. MGF ensures high-quality target features and suppresses speckle noise under SSL with SAR images. By taking advantage of these insights, SARatrX can learn high-quality target features from noisy SAR remote sensing images, as seen in Table II. In the next section, we analyze the diversity and scalability of SARatrX.
Figure 3: Discussions of single and multi-scale kernel settings for MGF. Here, the scale 1/2/3 assumes $r$ equal to 9/13/17, as the multi-scale contacts all scales. This multi-scale approach is more suitable than a single-scale technique for various targets in remote sensing images.
@@
Figure 4: Averaged attention distances for various attention heads (the x-axis is the attention head w.r.t layer number, and point colours represent different layers for better visualization) in the SSL models. Attention distance represents the range of a receptive field. We focused specifically on model architectures (Fig. (a) v.s. Fig. (b)), initialization weights (Fig. (a) v.s. Fig. (c)), and SSL signals (Fig. (d) v.s. Fig. (e)) to ensure diverse attention ranges for SAR target recognition, including the HiViT architecture, ImageNet weights, and SAR target features.
@@
Figure 5: The scalability of SARatrX for dataset size, parameters, and training epochs with linear probing performance. While our method benefited from these three attributes, it is important to note that excessive training epochs often led to overfitting, due to the dataset size and speckle noise.
@@
Visualization - Prior research [95] has shown that supervised pre-training and contrastive learning only model global information in higher layers, while MIM can model both local and global information. However, we observed that this effect was not only related to the chosen method but also to the data properties. Fig. 4 (a) demonstrates that ViT with MAE focuses on global information due to large SAR image scenes, which differs from MIM modeling properties. As such, HiViT includes various attention distances from 40 to 140 in different layers with its high spatial resolution and hierarchical structure, as seen in Fig. 4 (b). In addition, using ImageNet weights for initialization also solved this problem, as seen in Fig. 4 (c). Similarly, Fig. 4 (d) demonstrates that HOG features enhanced the noise interference, which limited feature diversity. MGF effectively extracts shape information from targets, focusing the model on diverse edge information in the lower layer. However, this approach removes textures and preserves edge information, which motivates higher layers to rely less on texture details and diminishes the attention range shown in Fig. 4 (e). Thus, we combined SSL-ImageNet &\&& SAR with MGF for two-step pre-training, as seen in Fig. 4 (f).
@@
Scaling experiment - Although MIM learns effectively and scales with data and model resources [96], a question arises as to whether our method can ensure scalability for MIM when dealing with noisy data, such as SAR. Fig. 5 presents the results of a scaling experiment from three perspectives: dataset size, parameters, and training epochs. Despite our pre-training set comprising 186,660 images, which is smaller than ImageNet-1K, we observed significant rising curve in downstream task performance with increasing data and parameter quantities in Figs. 5 (a) and (b). This result indicated that the foundational model could fully achieve its potential in SAR images by extracting high-quality features as guiding signals. However, as in [96], the model tended to overfit during extended training epochs when the pre-training set contained at least 100,000 images. In addition, SAR image noise and low resolution further aggravated the overfitting. Regardless, SARatrX outperformed our previous study (SAR-JEPA), which overfit at 400 epochs with 94,776 SAR images. Thus, there is a need to continue investigating new ways to ensure high-quality feature representations when extending SAR foundation models.
V Leveraging SARatrX for Recognition
@@
We have discussed different aspects of SARatrX, but there are several datasets and specialized models available for SAR ATR. Therefore, we compared our proposed SARatrX algorithm with other state-of-the-art techniques, such as supervised learning ($\rm{CS}^{n}$et [97] and PD [98]), semi-supervised learning (EUAPS [99]), and self-supervised learning (MSFA [21] and BIDFC [19]). We focused on SAR recognition tasks, including image classification and detection. More detailed settings [^4] We removed the test sets of downstream tasks from the pre-training set samples. are provided in Appendix B.
[…]
@@
SARatrX 2024 65.3 (+11.8) 76.5 (+20.3) 82.8 (+15.6)
[…]
@@
Classification task - Table IV details the performance of SARatrX for the MSTAR [18] dataset with standard operating conditions (SOCs) and extended operating conditions (EOCs). Notice that SSL (BIDFC and ours) and semi-supervised models (EUAPS) significantly outperformed other methods for small samples with additional unlabeled data. Our results also surpassed the previous best by large margins. For example, SOC 1-shot accuracy increased by 4.5% and EOCs 1-shot accuracy increased by 15.1% on average. demonstrating the value of FMs in an era of rapidly growing SAR data. In particular, SARatrX exhibited a robustness to the EOC setting of variable imaging conditions. This result indicated that the foundation model could learn stable features and relationships from diverse imaging conditions in a large number of samples.
Detection task - As illustrated in Table V, we reported a box mAP for SAR target detection with a horizontal bounding box for multi-category (SARDet-100K and OGSOD), ship (SSDD), and aircraft detection (SAR-Aircraft). SARatrX outperformed our previous MSFA by 0.8 points on SARDet-100K. MSFA also includes complex training processes and target features, which employ multi-stage training between RGB and SAR images with three different target features (HOG [83], Haar-like [84], WSTG [115]), including a detection pre-training step. SARatrX is thus simpler yet more effective for SAR images. Notably, SARatrX outperformed or offered comparable performance for multiple datasets, compared to several specifically designed detection methods shown in Table V.
Of course, our study is only a preliminary exploration of SSL for SAR image interpretation and more effective target features could be achieved in a data-knowledge dual-driven manner by further mining information on SAR imaging mechanisms and properties. Furthermore, given a larger dataset and computing power, the path of FMs will hopefully lead to generalized SAR interpretation include target recognition, scene classsification, semantic segmentation, and change detectionn, but this will require additional research.
VI Conclusion and Future Perspectives
@@
This study proposed SARatrX for SAR ATR and systematically investigated a foundation model framework. First, a pre-training dataset SARDet-180K was constructed from 14 open-source datasets, including various targets, scenes, and sensors. The foundation model’s pre-training backbone, SSL methods, and downstream tasks were then discussed in detail. Importantly, SARatrX demonstrated superior performance on different target recognition datasets, demonstrating the potential of FMs in this field. We believe that further research on SAR foundation models, including SARatrX, has the potential to generalize feature representations of SAR images and accelerate all-day, all-weather target recognition in Earth observations. However, foundation model research requires large data, which is problematic for SAR images, as they are expensive and require specific imaging equipment and algorithms. Privacy and security may also prevent the data from becoming open source. Therefore, we are particularly grateful to the publishers of open-source SAR target datasets. By making SARatrX publicly available, we aim to accelerate the progress of FMs in SAR target recognition by enabling researchers to use our dataset and code to design better methods or explore downstream applications.
@@
Although this work systematically investigated a foundation model framework, several limitations and challenges will require exploration in a future study. The SAR images were derived from open-source SAR datasets, and the targets primarily included vehicles, ships, aircraft, oil tanks, etc. Thus, collecting target samples from increasingly unlabeled imagery could further expand the amount of data to reach a million level and the range of downstream applications. In terms of model architecture, learning the dynamic contextual information on space-time is also an important issue. In addition to target shape features, various scenarios and target signatures for SAR self-supervised learning need to be explored. Finally, investigating expert knowledge with text used for multimodal interactions, and describing the relationship between targets and scenes, could further enhance the representation capabilities of the foundation model for SAR ATR.
@@
In conclusion, we have demonstrated the ability of SARatrX to adapt to diverse SAR target datasets, achieving high performance and generalizability in classification and detection tasks. By taking full advantage of the rapid growth of SAR images, this SSL-based foundation model opens the door to a generalized SAR target recognition.
Appendix A Implementation Details for Section 4 and SARatrX Experiments
@@
A-A Pre-traing dataset setting
The timeline of measured and simulated datasets used for SAR target recognition is illusate in Fig. 6, and we chose the typical open-source dataset as pre-training. As shown in Fig. 7, we collect data from open-source datasets based on our previous research [39, 21]. Now, our pre-training dataset contains 14 open-source SAR target datasets. Here are brief descriptions of each dataset’s targets, scenes, and sensors.
Figure 6: The timeline of measured and simulated datasets used for SAR target recognition (SAR magnitude images were processed in pseudo-color for better visualization). SAR target datasets have increased rapidly since 2020, greatly enriching the diversity of available targets, scenes, and sensors. The different colours represent the target categories included in the dataset. Target categories primarily include fine-grained vehicles, ships, and aircraft. A mixed sample implies the target exhibits more than one category, such as the combination of various vehicles, ships, and aircraft in IRIS-SAR. These scenes may also contain cities, harbors, airports, etc. The sensor devices offer different resolutions and bands for satellite or airborne platforms. However, the number of samples (No. samples) in most datasets is only a few thousand across several categories, due to high costs and the difficulty of annotation. This inspired us to propose SARatrX, with an integrated pre-training dataset SARDet-180K.
@@ A-A Pre-traing dataset setting
A-B Classification dataset for performance Test
We select three target classification datasets, including 25 fine-grained targets from vehicles, ships, aircraft, and others, to evaluate the comprehensive performance of SSL and the foundation model for SAR target recognition. The new SAR classification dataset named SAR-Target in Table VI.
A-C Hyperparameter settings
TABLE VI: Description of our SAR classification dataset, named SAR-VSA, which contains 25 fine-grained targets. # Train: Number of train samples. # Test: Number of test sample images.
@@ A-C Hyperparameter settings
TABLE VII: Pre-training setting.
@@
TABLE VIII: Classification settings (fine-tuning and linear probing).
@@
TABLE IX: Detailed classification results on MSTAR’s SOCs and EOCs. “-” refers to unknown. Bold indicates the best result, underline is the next best result.
@@
SARatrX 2024 65.3 76.5 76.3 83.4 82.8 85.0 92.1 97.0
TABLE X: Detailed detection results. The metrics are mAP, $\rm{mAP_{50}}$ (@@@@50), $\rm{mAP_{75}}$ (@@@@75), $\rm{mAP_{small}}$ (@@@@s), $\rm{mAP_{medium}}$ (@@@@m), and $\rm{mAP_{large}}$ (@@@@l). “-” refers to unknown. Bold indicates the best result, underline is the next best result.
@@
SARatrX 2024 58.7 86.1 (+5.7) 64.7 (+3.3) 20.0 57.2 49.7
Here are the detailed settings of our pre-training and downstream tasks as shown in Table VII and VIII.
Pre-training - Our default pre-training setting is in Table VII, and other hyperparameters of each method use the default settings from their papers and codes. Our pre-training is applied on 8 NVIDIA RTX3090 GPUs with 200 epochs and 800 batch sizes. Compared to the training settings of MAE, we add ColorJitter ($\rm{contrast}=0.5$) to increase data richness. Moreover, we modify the batch size and epoch according to 8 GPUs. It is worth noting that although MAE uses the normalized pixel value to enhance the feature representation in the visible spectral images, we find that the normalized pixel value cannot be used due to the SAR image noise and prevents the training loss from decreasing properly.
Classification setting - All models use the same training settings in downstream classification tasks. Table VIII gives the default setting. Our few-shot learning setting is based on the Dassl toolbox [121, 122] and is averaged over 10 random experiments. Since we focus on the small sample case of the downstream classification task, we use the linear probing method in MAE to finetune models and avoid overfitting.
Partial fine-tuning - Fig. 8 shows why we chose linear probing for few-shot evaluation, as HiViT overfits when many blocks are fine-tuned. Experimental results show that our SAR-KGPA consistently obtains better representations than MAE, and overfitting occurs later.
Appendix B Implementation Details for Section 5 and the Leveraging of SARatrX for Recognition Tasks
B-A Dataset description
We choose MSTAR, the most commonly used dataset in SAR target classification. SOCs are in similar image conditions, and the training set’s depression angle under SOC is 17°, while the test set is 15°. Ten categories of targets include BMP2, BRDM2, BTR60, BTR70, T62, T72, 2S1, D7, ZIL131, and ZSU234 as shown in Table VI. The EOCs setting is the imaging condition variations to test robustness. Existing methods have been saturated with different experimental settings of MSTAR [25], but few samples and depression angle variations remain challenging. We use SARDet-100K and OGSOD datasets, which have many test samples and categories, to evaluate the detection performance fully. The OGSOD comparison results are derived from the original article’s [22] single-modal approach using only SAR images. SSDD and SAR-Aircraft are ship and aircraft categories.
B-B Hyperparameter settings
Classification setting is follow Table VIII. The only difference is that we use partial fine-tuning for better performance, and the last 6 blocks are added to the fine-tuned.
B-C Detailed results
We provide detail classification and detection result on Table IX and X. Although the proposed method outperforms existing methods in mAP, there is still scope for improvement in some refined metrics.